The number of people who contracted a contagious disease t days after an epidemic started is approximated by the exponential equation:
Q(x) = 5000/1+1249e^-0.33t
Approximately how many people had contracted the disease after 15 days?

Answer by Cromlix(4381)

Q(x) = 5000/1+1249e^-0.33t
Q(x) = 5000/1 + 1249e^-0.33*15
Q(x) = 5000/1 + 1249e^-4.95
Q(x) = 507.75
Q(x) = 508 people.
Hope this helps.
:-)